# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

### v2

@@ IX Acknowledgments
Appendix A Particle in a Free Box
We consider simple models that clarify the treatment of momentum. The usual particle-in-a-box problem fixes the walls and therefore violates momentum conservation. Here the box is given finite mass and allowed to move freely, so that total momentum is exactly conserved.
Consider a one-dimensional rigid box of mass $M$ and length $L$ with center-of-mass coordinate $R$, containing a particle of mass $m$ with coordinate $r$ (Fig. 3).
Introduce total and relative coordinates
$\mathcal{R}=\frac{mr+MR}{m+M},\qquad\rho=r-R,$
with total mass $M_{T}=m+M$ and reduced mass
$\mu=\frac{mM}{m+M}.$
The hard-wall constraint is $|\rho|<L/2$.
The Hamiltonian separates exactly:
$H=-\frac{\hbar^{2}}{2M_{T}}\frac{\partial^{2}}{\partial\mathcal{R}^{2}}+\left[-\frac{\hbar^{2}}{2\mu}\frac{\partial^{2}}{\partial\rho^{2}}+V_{\rm hw}(\rho)\right].$
Total momentum is conserved:
$\hat{P}_{\rm tot}=-i\hbar\frac{\partial}{\partial\mathcal{R}}.$
Case A: Flat interior
The relative Hamiltonian is
$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm hw}(\rho).$
The normalized eigenfunctions are
$\phi_{n}(\rho)=\sqrt{\frac{2}{L}}\sin\!\left[\frac{n\pi}{L}\left(\rho+\frac{L}{2}\right)\right],\qquad n=1,2,\dots,$
with energies
$E_{n}=\frac{\hbar^{2}\pi^{2}n^{2}}{2\mu L^{2}}.$
Any internal state may be written
$\psi(\rho,t)=\sum_{n=1}^{\infty}c_{n}\,\phi_{n}(\rho)\,e^{-iE_{n}t/\hbar}.$
A Gaussian superposition centered at $\rho_{0}$ with mean wavenumber $k_{0}$,
$c_{n}=\mathcal{N}\exp\!\left[-\frac{(n-n_{0})^{2}}{4\sigma_{n}^{2}}\right]\exp\!\left[-ik_{n}\rho_{0}\right],\qquad k_{n}=\frac{n\pi}{L},$
produces a localized packet
$\psi(\rho,0)\approx A(\rho-\rho_{0})e^{ik_{0}(\rho-\rho_{0})}.$
Although each $\phi_{n}$ carries no relative momentum, the particle and box momenta may be nonzero individually, while $P_{\rm tot}$ remains fixed. At a wall collision both reverse sign, preserving total momentum.
Case B: Periodic interior
Let the interior potential be periodic relative to the box:
$V_{\rm per}(\rho)=V_{\rm per}(\rho+a).$
Then
$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm per}(\rho)+V_{\rm hw}(\rho).$
Eigenfunctions take Bloch standing-wave form
$\varphi_{bn}(\rho)=\sqrt{\frac{2}{L}}u_{bn}(\rho)\sin\!\left(\frac{n\pi}{L}(\rho+L/2)\right),$
with
$u_{bn}(\rho+a)=u_{bn}(\rho).$
Wavepackets are superpositions
$\psi(\rho,0)=\sum_{n=1}^{\infty}A_{n}\varphi_{bn}(\rho).$
Shared pseudomomentum
Because $\rho=r-R$,
$e^{iq\rho}=e^{iqr}e^{-iqR}.$
Internal pseudomomentum $\hbar q$ is therefore shared between particle and box. For fixed total momentum $P_{0}$,
$p_{r}=\frac{m}{m+M}P_{0}+\hbar q,\qquad P_{R}=\frac{M}{m+M}P_{0}-\hbar q,$
so $p_{r}+P_{R}=P_{0}$.
Cases C and D: Dynamical periodic background
Let the periodic background be a dynamical elastic chain with atomic coordinates $R_{j}$:
@@ A Gaussian superposition centered at $\rho_{0}$ with mean wavenumber $k_{0}$,
With density–density coupling
@@
Thus total momentum
$P_{\rm tot}=p_{r}+\sum_{j}p_{j}$
is exactly conserved. Diagonalizing the lattice,
$R_{j}=R_{\rm lat}+\sum_{q}u_{q}e^{iqja},$
reveals the zero mode $R_{\rm lat}$. The density operator contains the factor
$e^{-iqR_{\rm lat}},$
showing explicitly that momentum $\hbar q$ transferred to the particle is balanced by recoil of the lattice center of mass. Internal phonons redistribute momentum but are not required for momentum conservation.
Appendix B Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
We consider a finite one-dimensional tight-binding chain of $N$ identical atoms, isolated in space, with nearest-neighbor springs and an electron moving in a basis of localized orbitals. The chain possesses an exact translational zero mode (center-of-mass motion). An electron can therefore exchange momentum with the lattice as a whole without necessarily exciting internal phonon modes. This opens a phonon-diagonal (internally elastic) backscattering channel for an electronic wavepacket, with the required momentum reversal balanced by center-of-mass recoil.
Hamiltonian and separation of the CM mode
Write the atomic coordinates as
$x_{n}=R+na+u_{n},\qquad\sum_{n=1}^{N}u_{n}=0,$ (41)
where $R$ is the center-of-mass coordinate, $a$ the equilibrium lattice spacing, and $\{u_{n}\}$ the internal displacements. The lattice Hamiltonian separates into CM and internal parts,
$H_{\rm lat}=\frac{P_{R}^{2}}{2M_{\rm tot}}+\sum_{n=1}^{N}\frac{p_{n}^{2}}{2M}+\frac{K}{2}\sum_{n=1}^{N-1}(u_{n+1}-u_{n})^{2},$ (42)
with $M_{\rm tot}=NM$ and $\sum_{n}p_{n}=0$.
The electron is described by a nearest-neighbor tight-binding Hamiltonian,
$H_{e}=-\sum_{n=1}^{N-1}t_{n}\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right),$ (43)
with bond-length-dependent hopping
$t_{n}\equiv t(x_{n+1}-x_{n})\simeq t_{0}-\alpha(u_{n+1}-u_{n}).$ (44)
Thus the electron–lattice coupling depends only on internal coordinates $u_{n+1}-u_{n}$; the CM coordinate $R$ drops out exactly.
Phonon-diagonal backscattering via CM recoil
Because the full electron+lattice Hamiltonian is translationally invariant, the total momentum is exactly conserved. Since the band energy satisfies $\varepsilon(k)=\varepsilon(-k)$, electronic reversal $k\to-k$ does not change the electronic energy.
Consider an incoming wavepacket narrowly peaked at +k+k, an internal lattice eigenstate $|\nu\rangle$, and a CM momentum eigenstate $|P_{R}\rangle$. A phonon-diagonal backscattering channel has the form
$|k\rangle\otimes|\nu\rangle\otimes|P_{R}\rangle\longrightarrow|-k\rangle\otimes|\nu\rangle\otimes|P_{R}+2\hbar k\rangle,$ (45)
i.e. the internal state is unchanged while the CM absorbs the recoil $2\hbar k$.
The associated change in CM kinetic energy is
$\Delta E_{\rm CM}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{4\hbar k\,P_{R}+4\hbar^{2}k^{2}}{2M_{\rm tot}}.$ (46)
Exact energy conservation with strictly unchanged internal state requires $\Delta E_{\rm CM}=0$, i.e.
$P_{R}=-\hbar k.$ (47)
For generic $P_{R}$ there is a recoil-energy mismatch
$\Delta E_{\rm recoil}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{2\hbar k\,(P_{R}+\hbar k)}{M_{\rm tot}}.$ (48)
However, for a macroscopic chain $M_{\rm tot}\propto N$ so $|\Delta E_{\rm recoil}|\sim 1/N$ becomes parametrically small. For a wavepacket of spatial extent $L$ with intrinsic energy width $\Delta E_{\rm pkt}\sim\hbar v_{g}/L$, or for scattering occurring over a finite time $\tau_{\rm scatt}$ with uncertainty $\hbar/\tau_{\rm scatt}$, the process is effectively elastic whenever $\Delta E_{\rm recoil}$ is below the available energy resolution. In a number-state phonon description, any residual mismatch may be accommodated by very soft internal modes, but the momentum reversal $k\to-k$ is carried by CM recoil; internal excitations, if present, serve primarily as an energetically low-cost bookkeeping channel.
Role of finite temperature
At finite temperature $T$, the CM momentum is thermally distributed,
$f(P_{R})\propto\exp\!\left(-\frac{P_{R}^{2}}{2M_{\rm tot}k_{B}T}\right),$ (49)
with width $\sigma_{P}=\sqrt{M_{\rm tot}k_{B}T}$. The resonant value $P_{R}=-\hbar k$ therefore occurs with nonzero probability density for any $T>0$, so strictly phonon-diagonal backscattering events exist at finite temperature in an isolated chain.
Bulk and boundary backscattering
The CM-recoil channel does not require the electron to reach the ends: it can occur in the bulk from the static component of thermal distortions (phonon-diagonal scattering). The SSH coupling
$H_{\rm int}=\alpha\sum_{n=1}^{N-1}(u_{n+1}-u_{n})\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right)$
connects $|k\rangle$ and $|-k\rangle$ through the $2k$ Fourier component of the bond-distortion field. CM recoil is enforced by the CM translation factor. In particular, the CM shift operator produces
$\langle P_{R}^{\prime}|e^{-i(2k)R}|P_{R}\rangle\propto\delta(P_{R}^{\prime}-P_{R}-2\hbar k),$ (50)
so momentum conservation is implemented by a delta function (or Kronecker delta in a finite quantization volume), not by an inner product between distinct CM momentum eigenstates.
Free boundaries can also reflect a wavepacket once it reaches an end. Such reflections reverse $k$ and are likewise accompanied by CM recoil. Boundaries provide a location for reflection but do not constitute a momentum sink; both bulk (distortion-mediated) and boundary reflections obey the same conservation principle enforced by the CM zero mode.
Other scattering channels
In addition to the phonon-diagonal CM-recoil channel, inelastic and hybrid processes are allowed.
Single-phonon emission and absorption
If momentum is transferred to internal modes, the internal lattice state changes:
$|k\rangle\otimes|\{n_{q}\}\rangle\to|k^{\prime}\rangle\otimes|\{n_{q}^{\prime}\}\rangle,\qquad n_{q}^{\prime}=n_{q}\pm 1.$
Energy conservation requires
$\varepsilon(k^{\prime})=\varepsilon(k)\pm\hbar\omega_{q}+\Delta E_{\rm CM}.$
A Golden-Rule rate has the form
$W_{k\to k^{\prime};q}=\frac{2\pi}{\hbar}|M_{k,k^{\prime};q}|^{2}\,\delta\!\left(\varepsilon(k^{\prime})-\varepsilon(k)\mp\hbar\omega_{q}-\Delta E_{\rm CM}\right),$
with matrix elements scaling as
$M_{k,k^{\prime};q}\sim\alpha\sqrt{\frac{\hbar}{2M\omega_{q}}}\,F_{k,k^{\prime};q}.$
Multi-phonon and hybrid CM–phonon processes
Higher-order multi-phonon processes and hybrid processes involving both CM recoil and internal phonon exchange relax the strict resonance condition (47) at the cost of reduced matrix elements.
Kinematic summary
  • Phonon-diagonal channel: $k\to-k$ with unchanged internal state is allowed by translational invariance and proceeds via CM recoil; it is strictly elastic only on resonance (47) and effectively elastic for large $N$.
  • Inelastic channels: internal phonon emission/absorption become important when the phonon spectrum permits energy matching.
  • Finite size: discrete internal spectra constrain inelastic channels and can enhance the relative importance of CM-mediated phonon-diagonal scattering.
Practical consequences
  • In wavepacket simulations, phonon-diagonal backscattering is identified by momentum reversal with unchanged internal state and a compensating shift of $P_{R}$ by $2\hbar k$.
  • Bulk and boundary reflections differ only in spatial location; both satisfy the same momentum conservation enforced by the CM zero mode.
  • Neglecting the CM zero mode (implicit clamping) qualitatively changes the available scattering channels.
@@ Appendix C Whither Umklapp?
The traditional role of Umklapp for finite resistivity in an infinite, perfect crystal arises from freezing the background and projecting out its pseudomomentum. Once the background is treated as dynamical, with an explicit translational degree of freedom, momentum relaxation can occur through elastic pseudomomentum exchange between electron and lattice. Phonons then represent one possible basis for describing background dynamics rather than the exclusive channel for momentum transfer.
This perspective clarifies long-standing ambiguities surrounding Umklapp and explains why momentum relaxation is observed experimentally even in regimes where Umklapp processes are expected to be suppressed. It also underscores that the association of Umklapp with dissipation and entropy production is not intrinsic, but follows from additional assumptions about the background.
The expectation that Umklapp processes are necessary for finite resistivity in a perfect crystal has deep roots in transport theory. Peierls first distinguished between normal and Umklapp processes in his analysis of phonon thermal conductivity, showing that only the latter can relax total crystal momentum in a closed system. This reasoning was later extended to electrons by Bloch, Peierls, Ziman, and others, and appears explicitly in standard texts (e.g. Peierls, Quantum Theory of Solids; Ziman, Electrons and Phonons). If the Fermi surface lies well inside the Brillouin zone, or at low temperature where large-wavevector phonons are thermally suppressed, Umklapp processes are expected to be exponentially rare. Under these conditions conventional theory predicts extremely long momentum lifetimes in a clean crystal.
The present formulation revisits this conclusion by restoring the dynamical translational degree of freedom of the background. Momentum exchange between subsystems does not require phonon creation, and the strict identification of Umklapp with momentum relaxation relies on additional assumptions about background clamping or environmental coupling.
Umklapp processes are best understood as a classification of crystal-momentum conservation rather than as the fundamental cause of momentum relaxation. Their elevated role in conventional transport theory stems from treating the lattice background as inert. Recognizing the background as a dynamical object carrying its own pseudomomentum restores a more symmetric and physically complete description in which elastic momentum exchange is allowed and Umklapp is no longer indispensable.

### v3

@@ XI Acknowledgments
Appendix A Bloch waves form from elastic scattering of electrons in a lattice
By way or reinforcing the conclusion that elastic internal electron scattering must be very common in pure metals, we raise a point about the rather violent elastic scattering that is taking place, cl oacked by the symmetry of Bloch waves. The formation of Bloch waves in a periodic crystal may be viewed as the result of coherent multiple elastic scattering from the atoms of the lattice[20]. Band formation itself is therefore an elastic scattering phenomenon. One might then suppose that Bloch waves are somehow protected from further elastic deflection by phase coherence. However, this inference is incorrect, as illustrated in Fig. 5.
We extend this insight slighty and suppose some atom were temporarily slightly out of place due to thermal acoustic waves. The atoms would individually scatter as before, but now the phase locking is imperfect, and the electron can deflect. There is no new “phonon needed” physics that sets in due to truly minuscule thermal displacements. If the electron thus deflects, momentum is exchanged with the lattice without phonon excitation or energy dissipation.
Appendix B Tracking Momentum, and the Effects of Zero Mode Removal
Figure 6 helps illuminate the choices that can be made in treating electron-phonon interaction. It helps to underline that there is a great difference between treating the background as a dynamical body (with a translational zero mode) or as frozen, without recoil. Frozen is an impossibility in the real world, total mometum is always conserved, although this a self-consistent model. The question is, does that model lead to good real-world predictions, or does it mislead? Retaining the zero mode total momentum is not a matter of interpretation; it fixes the form of the wavefunction and alters the physical predictions in important ways.
Translation invariance of the whole requires that the Hamiltonian depends only on the center of mass and coordinate differences, so eigenstates separate into a plane wave for the center of mass coordinate and an internal state that depends only on relative coordinates.
B.1 Uplifting the particle in a box to the real world.
We therefore free the textbook particle in a box from its traditional and physically impossible bondage in case A.: considering a“free floating box” of mass $M$ containing a particle of mass $m$, with particle coordinate $r$ and box center-of-mass coordinate $R$, introduce
$\mathcal{R}=\frac{mr+MR}{m+M},\qquad$
Then a complete set of eigenstates (delta-normalized in $P$) may be written as
$\Psi_{P,n}(r,R)=\frac{1}{\sqrt{2\pi\hbar}}\,e^{iP\mathcal{R}/\hbar}\,\phi_{n}(r-R).$ (110)
Equation (110) is the operative content of keeping the zero mode: the internal structure is literally a function of rRr-R. The conserved total mechanical momentum is
$\hat{P}_{\rm tot}=-i\hbar\frac{\partial}{\partial\mathcal{R}},\qquad[H,\hat{P}_{\rm tot}]=0.$
Any internal state may be expanded in Fourier components on |rR|<L/2|r-R|<L/2,
$\phi_{n}(r-R)=\sum_{q}a_{nq}\,e^{iq(r-R)},$
and each component carries the shared phase
$e^{iq(r-R)}=e^{iqr}e^{-iqR}.$ (111)
Thus, in the plane-wave decomposition of the relative coordinate, momentum exchange is enforced at the level of the wavefunction: each Fourier component correlates $\hbar q$ on the particle with $\hbar q$ on the background. Phonons and other internal excitations may redistribute momentum within the background.
B.2 Box is corrugated
Suppose the box is as above but now contains a rigid, corrugated periodic potential (Fig. 6, B). Before the corrugation, the particle and the box each carry mechanical momentum, and only the total $P_{\rm tot}=p+P$ is conserved. In collisions, momentum is exchanged between the two in equal and opposite amounts, exactly as in free space.
The introduction of a periodic potential does not eliminate this structure, but reorganizes it. Continuous translational symmetry is reduced to discrete translational symmetry, and the corresponding conserved quantities become pseudomomenta defined modulo reciprocal lattice vectors. Importantly, the two-body structure of momentum exchange is preserved.
The particle now carries a pseudomomentum $\hbar k$, while the background box carries a corresponding pseudomomentum. These two pseudomomenta are exchanged in equal and opposite amounts during scattering, just as the mechanical momenta were exchanged before the corrugation was introduced. Their sum remains fixed and corresponds to the conserved total momentum of the isolated system.
Thus the corrugation does not replace two mechanical momenta with a single pseudomomentum; rather, it produces two pseudomomenta, one for the particle and one for the background. The center-of-mass degree of freedom of the box continues to carry the total mechanical momentum, while the periodic structure introduces a new, discrete labeling of the same underlying exchange process.
The situation depicted in Fig. 6, B therefore retains the essential physics of momentum exchange: the lattice remains a dynamical participant, even though its role is partially obscured by the use of pseudomomentum.
B.3 Elastic Chain in a Box, Free CM (phonons plus the zero mode)
Now take the background to be a dynamical chain (pink atoms) of total mass $M_{\rm lat}$ inside the box, with atomic coordinates $R_{j}$ and a center-of-mass coordinate $R_{\rm lat}$ (Fig. 6, C). A minimal model is
@@ B.3 Elastic Chain in a Box, Free CM (phonons plus the zero mode)
With density-density coupling
@@
Consequently the conserved quantity is the true total mechanical momentum
$P_{\rm tot}=P_{e}+P_{\rm lat},\qquad[H,P_{\rm tot}]=0.$
Momentum exchange is enforced directly through the relative coordinate. Fourier transforming the lattice density yields
$\rho_{\rm lat}(q)=e^{-iqR_{\rm lat}}\rho^{\rm int}_{q},\qquad H_{\rm int}\propto e^{iq(r-R_{\rm lat})}.$
The phase factor $e^{iq(r-R_{\rm lat})}$ encodes the fundamental constraint: momentum $\hbar q$ transferred to the electron is accompanied by $\hbar q$ carried by the lattice center of mass. Internal degrees of freedom describe only relative motion and carry no net mechanical momentum.
The translational zero mode $R_{\rm lat}$ is therefore the unique carrier of the compensating momentum. Internal modes (phonons) may change the energy of the system, but they are not required to enforce momentum conservation.
Deleting the zero mode is not an innocuous simplification. It removes the only degree of freedom capable of carrying recoil and eliminates the shared-coordinate structure $\phi(r-R)$ from the Hilbert space. The resulting theory no longer represents true mechanical momentum conservation; instead it replaces it with pseudomomentum bookkeeping within a reduced space. Momentum exchange is then no longer enforced through the relative coordinate rRr-R, but appears only as a selection rule requiring phonon pseudomomentum to balance electron pseudomomentum (up to Umklapp).
Thus the apparent kinematic necessity of phonon production does not arise from fundamental principles. It is imposed by amputating the translational degree of freedom that would otherwise carry the compensating momentum.
B.4 Free Chain (no box), Free CM (translation invariance made manifest)
Removing the box entirely leaves a floating chain. This is simply Fig. 6, C with boundary constraints eliminated. The lesson is unchanged but becomes explicit: the conserved momentum is the generator of rigid translation of all coordinates.
Any description that retains translational invariance necessarily contains the shared phase $e^{iq(r-R_{\rm lat})}$, since it is the Fourier representation of the relative coordinate. This phase structure is the direct expression of momentum conservation in the full system.
B.5 Removing the Zero Mode
The conventional electron-phonon kinematics is obtained by a construction that is the Hilbert-space analog of “nailing the box down,” but without introducing an external dynamical reservoir. The translational zero mode is removed at the outset. One expands the lattice displacement field in normal modes and omits the $q=0$ component. The resulting description retains only internal degrees of freedom.
With the zero mode removed, continuous translation symmetry is no longer present, and the conserved quantity is not the true mechanical momentum. What remains is discrete lattice translation symmetry, with conserved crystal pseudomomentum
$K_{\rm tot}=K_{e}+K_{\rm ph}\quad(\mathrm{mod}\;G),\qquad K_{\rm ph}=\sum_{q}\hbar q\,a_{q}^{\dagger}a_{q},$
and the corresponding selection rule
$\Delta K_{e}+\Delta K_{\rm ph}=Gm.$
In this representation, momentum transfer is confined to phonons: changes in the electron pseudomomentum must be balanced entirely by phonon pseudomomentum, up to Umklapp. The lattice center-of-mass degree of freedom, which would otherwise provide a direct channel for momentum exchange through rigid recoil, is absent.
The apparent kinematic necessity of phonon emission or absorption is therefore not a physical result. It is a consequence of having eliminated the translational degree of freedom that would otherwise carry the compensating momentum. It is not a choice one normally would wish to make, because it changes the predicted physical processes. It is surprising perhaps, but it does. The “thermodynamic limit” does not protect us here, and if it seems that it does, the limit $N\to\infty$ has not been taken properly. Jumping to $N=\infty$ is not taking the limit of course.
@@ Appendix C Whither Umklapp?
The thermodynamic limit and Umklapp processes are often invoked to account for momentum relaxation in perfect crystals. These mechanisms are valid within the conventional formulation, but their apparent necessity reflects the prior removal of the lattice center-of-mass (zero-mode) degree of freedom.
In the standard phonon description, momentum conservation is imposed only at the level of crystal pseudomomentum,
$K_{\rm tot}=K_{e}+K_{\rm ph}\quad(\mathrm{mod}\;G),$
and electronic momentum can relax only through phonon processes, supplemented by Umklapp when reciprocal lattice vectors are invoked. This framework is kinematically restrictive: the phonon system tracks momentum only modulo $G$, and the available channels are limited to discrete reciprocal lattice transfers.
When the zero mode is retained, this restriction is lifted. The lattice background can recoil and absorb arbitrary mechanical momentum, and the scattering problem is no longer confined to Umklapp processes. Instead, momentum transfer occurs within a continuous manifold of elastic and quasielastic processes involving the full dynamical background.
To make this explicit, we separate a single electron (the foreground) from the rest of the system (the background), which includes all ions and all other electrons. The total mechanical momentum is
$\hat{P}_{\rm tot}=\hat{p}+\hat{P}_{\rm bg},\qquad\hat{P}_{\rm bg}=\sum_{\alpha\in\mathrm{bg}}\hat{P}_{\alpha}.$ (112)
The operator $\hat{P}_{\rm bg}$ generates rigid translation of the background and constitutes its center-of-mass (zero-mode) momentum. Internal phonon modes describe only relative motion and do not furnish an independent contribution to $\hat{P}_{\rm bg}$. Thus any change in the electron’s mechanical momentum must be balanced by the background zero mode.
This makes clear what is implicit in Umklapp. When a reciprocal lattice vector $G$ appears in a scattering process, the associated momentum $\hbar G$ cannot be attributed to the electron alone, nor is it carried by internal phonon modes. It must be carried by the lattice as a whole. In the conventional formulation, however, the zero mode that would carry this momentum has been removed, and the role of recoil is replaced by bookkeeping modulo $G$.
The traditional role assigned to Umklapp in producing finite resistivity therefore arises within a restricted description in which the background is effectively clamped. Once the background is treated dynamically, electronic momentum need not decay through Umklapp. Instead, it can be transferred elastically to the lattice as a whole.
Umklapp remains a useful classification of pseudomomentum conservation modulo reciprocal lattice vectors. It is not, however, the unique or even generic mechanism for momentum relaxation in a real crystal, where the background is free to recoil.
Appendix D The landscape seen by an electron
At temperatures of order $T\sim 100\,$K, thermally excited acoustic distortions in a typical metal generate substantial deformation-potential landscapes on nanometer length scales. Using standard values for the deformation potential ($g\sim 5$$15\,\mathrm{eV}$), sound velocity ($v_{s}\sim 3$$5\times 10^{3}\,\mathrm{m/s}$), and the characteristic thermal phonon wavevector $q_{T}\sim k_{B}T/\hbar v_{s}$, one finds order-of-magnitude estimates for the resulting deformation-potential field gradients of
[…]